# Every finite set of natural numbers is realizable as algebraic periods of a Morse–Smale diffeomorphism

### v1

@@ 3. Proof of main theorem
In the forthcoming considerations we will use the notion of $k$-periodic expansion of an arithmetic function $\psi:\mathbb{N}\to\mathbb{C}$ introduced in [Mar-NP], exposed widely in [Je-Ma], and used in [GLM], [Gr-L-NP]. In this language an arithmetic function $\psi$ is represented as a series of elementary periodic functions
@@
where the coefficients $a_{k}(\psi)\in\mathbb{C}$ are given by the Möbius inversion formula as in (1) and
@@
is the sum of $n$-powers of all roots of unity of degree $k$ and is called the elementary $k$-periodic function. Conceptually, the $k$-periodic expansion is a discrete correspondent of the Fourier expansion, arithmetically it is linked with the Ramanujan sums, and its coefficients are used to study the behaviour of the number sequences which are related to the periodic points (see [EPPW]). In our context, it is just a convenient language to handle the sequence we are studying.
@@
    There exists a closed orientable surface $S_{\mathrm{g}}$ and an orientation-preserving homeomorphism $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ and the genus $\mathrm{g}$ of $S_{\mathrm{g}}$ is equal to
@@
    There exists a closed orientable surface $S_{\mathrm{g}}$ and an orientation-reversing homeomorphism $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ if and only if ${\mathcal{A}}\subset 2{\mathbb{N}}$. Moreover, the genus $\mathrm{g}$ of $S_{\mathrm{g}}$ is equal to
@@
    There exists a closed non-orientable surface $N_{\mathrm{g}}$ and a homeomorphism $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ and the genus $\mathrm{g}$ of $N_{\mathrm{g}}$ is equal to
@@
First, let us describe the following construction. Having a finite set ${\mathcal{A}}^{\prime}\subset{\mathbb{N}}$ and an assignment $\tau\colon{\mathcal{A}}^{\prime}\to{\mathbb{N}}$, for each $n\in{\mathcal{A}}^{\prime}$ consider a surface $\Sigma_{n}$, which is an orientable surface $S_{\tau(n),2}$ in the orientable case or a non-orientable surface $N_{\tau(n),2}$ in the non-orientable case. On each $\Sigma_{n}$ we take a periodic homeomorphism $f_{n}$ of order $\tau(n)$ which cyclically permutes its $1$-handles and is a rotation on each boundary component (see Figure 1). More specifically, for an illustration, start with a cylinder $S^{1}\times[-1,1]$ with a rotation by $1/\tau(n)$ of a full angle on each circle $S^{1}\times\{t\}$ and remove $\tau(n)$ open disjoint discs $D_{1},\ldots,D_{\tau(n)}$ in such a way that the rotation cyclically permutes the discs. Now glue $\tau(n)$ copies of $S_{1,1}$ (or $N_{1,1}$) along permuted boundary components and extend the rotation to a periodic homeomorphism of order $\tau(n)$ that cyclically permutes the attached handles.
Next, having surfaces $\Sigma_{n_{1}},\ldots,\Sigma_{n_{k}}$ let us join $\Sigma_{n_{i}}$ and $\Sigma_{n_{i+1}}$ for $11\leq i\leq k-$ 1 by identifying their boundary components with end circles of a cylinder $C_{i}=S^{1}\times[-1,1]$. By the construction, we get a surface $S$ with two boundary components and define $f$ to be equal to $f_{n_{i}}$ on $\Sigma_{n_{i}}$, and on each joining cylinder it is given by a homotopy between two rotations on end circles.
@@
Figure 1.
(1) In the case of an orientable surface and an orientation-preserving homeomorphism, take ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{1\}$ if $1\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{1\}$ if $1\in{\mathcal{A}}$, with an assignment $\tau(n)=n$ for $n\in{\mathcal{A}}^{\prime}$. The above procedure provides a surface $S$ with an orientation-preserving homeomorphism $f$. Let us attach two discs to $S$ to get a closed orientable surface $\tilde{S}$, and $f$ extends on $\tilde{S}$, being an appropriate rotation on each of two discs. Taking one circle $S^{1}\times\{0\}$ for each cylinder $C_{i}$ we get a system of simple closed curves preserved by $f$ such that $f$ is periodic on the complement of its open tubular neighbourhood $\bigcup{\rm Int}C_{i}$, so $f$ is of algebraically finite type. Moreover, let $B_{n}=\{a^{n}_{1},\ldots,a^{n}_{n},b^{n}_{1},\ldots,b^{n}_{n}\}$ consist of circles in $S_{n}$ such that $\bigcup_{n\in{\mathcal{A}}^{\prime}}B_{n}$ forms the standard (symplectic) basis for $\tilde{S}$. Then the matrix of $H_{1}(f)$ on $H_{1}(\tilde{S})$ is a block diagonal matrix whose blocks correspond to $B_{n}$’s and since $f(a^{n}_{j})=f_{n}(a^{n}_{j})=a^{n}_{(j+1)\mod n}$ and similarly $f(b^{n}_{j})=b^{n}_{(j+1)\mod n}$, the $n$th block is the direct sum of two permutation matrices for a cycle of length $n$:
@@
and so it is not difficult to observe by the definition of ${\rm reg}_{k}$ that: $\mathrm{t}r(H_{1}(f^{l}))=\sum_{n\in{\mathcal{A}}^{\prime}}2{\rm reg}_{n}(l)$. By the definition of ${\mathcal{A}}^{\prime}$,
@@
where $a_{n}(f)=-2$ for $n\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$ and $a_{n}(f)=0$ for $n\notin{\mathcal{A}}$. Thus ${\mathcal{A}P}(f)={\mathcal{A}}$.
(2) For the orientation-reversing case, we will show later that there are no odd algebraic periods (see Theorem 3.2). Assume that ${\mathcal{A}}\subset 2{\mathbb{N}}$ and take ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{2\}$ if $2\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{2\}$ if $2\in{\mathcal{A}}$ with an assignment $\tau(n)=n$ if $4$ divides $n$ and $\tau(n)=n/2$ otherwise. Again, the procedure provides an orientable surface $S$, and we attach a disc to one of its two boundary components (name the resulting surface also $S$). We take its double $D(S)$, i.e. consider a copy $S^{\prime}$ of $S$ consisting of surfaces $\Sigma^{\prime}_{n}$ and cylinders $C^{\prime}_{i}$, and join $S$ and $S^{\prime}$ by a new cylinder $C=S^{1}\times[-1,1]$ along their boundary components. We have defined $f$ on $S$ and $S^{\prime}$, and on the new cylinder, let $f_{|S^{1}\times\{t\}}$ be the same as $f_{|{S^{1}\times\{\pm 1\}}}$.
Now, let $j\colon D(S)\to D(S)$ be an orientation-reversing involution mapping points of $S$ into corresponding points of $S^{\prime}$, e.g. it can be seen as the reflection through the plane intersecting $D(S)$ in $S^{1}\times\{0\}\subset C$ and separating symmetrically $S$ and $S^{\prime}$, see Figure 1 (2). The desired self-homeomorphism of $D(S)$ is $f^{\prime}=f\circ j$. Indeed, cutting $D(S)$ along circles from connecting cylinders, we get the decomposition into surfaces $\Sigma_{n}$ and $\Sigma^{\prime}_{n}$ such that $f^{\prime}$ maps $\Sigma_{n}$ homeomorphically onto $\Sigma^{\prime}_{n}$ and conversely. Thus $(f^{\prime})^{2}$ is a self homeomorphism of each piece $\Sigma_{n}$ or $\Sigma^{\prime}_{n}$ on which one can check it is $n/2$-periodic. Hence $f^{\prime}$ is of algebraically finite type.
Moreover, it is clear that $(f^{\prime})^{2}(a^{n}_{j})=a^{n}_{(j+2)\mod\tau(n)}$, $(f^{\prime})^{2}(b^{n}_{j})=b^{n}_{(j+2)\mod\tau(n)}$ on $\Sigma_{n}$ and similarly for $\Sigma^{\prime}_{n}$. If $4$ divides $n$, then $\tau(n)=n$ and so $(f^{\prime})^{2}_{|\Sigma_{n}}$ has order $n/2$, thus the $n$th block of the matrix of $H_{1}(f^{\prime})$ is the direct sum of four permutation matrices for a cycle of length $n$. However, if $n$ is not divisible by $4$, then $\tau(n)=n/2$ is odd and $(f^{\prime})^{2}_{|\Sigma_{n}}$ still has order $\tau(n)=n/2$. Therefore the characteristic polynomial of $f^{\prime}$ is equal to
@@
(3) Analogously we construct a desired self-homeomorphism of a non-orientable surface. If ${\mathcal{A}}=\{1\}$, then just take ${\rm id}_{{\mathbb{R}}\!P^{2}}$. Otherwise, consider ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{1\}$ if $1\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{1\}$ if $1\in{\mathcal{A}}$ with an assignment $\tau(n)=n$ for $n\neq 1$ and $\tau(1)=2$ if $1\in{\mathcal{A}}^{\prime}$. Moreover, in the preliminary construction for the non-orientable case if $1\in{\mathcal{A}}^{\prime}$ we take $f_{1}={\rm id}$ on $\Sigma_{1}$ which is of genus $\tau(1)=2$. We get a non-orientable surface $S$ with two boundary components to which we attach discs as previously. Similarly as in (1) we have a self-homeomorphism $f$ of $S$ of algebraically finite type, which is $n$-periodic on each $\Sigma_{n}$, except $n=1$.
Let $B_{n}=\{a^{n}_{1},\ldots,a^{n}_{\tau(n)}\}$ be a set of circles in $\Sigma_{n}$ which are cyclically permuted by $f$ and such that elements of $\bigcup B_{n}$ represent standard generators of $H_{1}(S)\cong{\mathbb{Z}}^{g-1}\oplus{\mathbb{Z}}_{/2}$ with the relation $02\sum_{i,n}a^{n}_{i}=$ 0, where $\mathrm{g}$ is the genus of $S$. The trace of $H_{1}(f)$ is computed as the trace of $\widetilde{H_{1}(f)}$ on $H_{1}(S)/{\rm T}\cong{\mathbb{Z}}^{\mathrm{g}-1}$, which is induced by $H_{1}(f)$ after dividing by the torsion part ${\rm T}$ of $H_{1}(S)$. Thus $H_{1}(S)/{\rm T}\cong\left(\bigoplus_{i,n}{\mathbb{Z}}a^{n}_{i}\right)/\left<\sum_{i,n}a^{n}_{i}\right>$ has a basis $\left(\bigcup B_{n}\right)\setminus\{a\}$ for some fixed $a\in B_{n_{0}}$ and $a$ is represented as the vector $(-1,\ldots,-1)$ in this basis. Hence the matrix of $\widetilde{H_{1}(f)}$ is a block matrix if $n_{0}=1\in{\mathcal{A}}^{\prime}$ or almost a block matrix if $n_{0}\neq 1$, i.e. the structure of a block matrix is disturbed only in the column which corresponds to a basis element $a^{n_{0}}_{i}$ such that $f(a^{n_{0}}_{i})=a$ (if we order the basis in such a way that $a^{n_{0}}_{i}$ is the last element, then the matrix is upper block triangular). However, in the second case the characteristic polynomial of $\widetilde{H_{1}(f)}$ is still the product of characteristic polynomials of permutation matrices for cycles of length $n$ for $n\neq n_{0}$, the identity matrix of dimension $2$ if $1\in{\mathcal{A}}^{\prime}$ and the characteristic polynomial of
@@
which is the companion matrix of the polynomial $x^{n_{0}-1}+\ldots x+1=(x^{n_{0}}-1)/(x-1)$. Since the trace of the companion matrix of a polynomial is equal to the sum of its roots, by the formula (3), the trace of $l$-power of the matrix (5) is equal to ${\rm reg}_{n_{0}}(l)-{\rm reg}_{1}(l)$. Thus in both the cases we get:
@@
where $a_{n}(f)=-1$ for $n\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$ and $a_{n}(f)=0$ for $n\notin{\mathcal{A}}$. Hence ${\mathcal{A}P}(f)={\mathcal{A}}$.
To complete the proof, we are left with showing that ${\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)=\emptyset$ if $f$ is orientation-reversing, because we have shown only this realization for ${\mathcal{A}P}(f)\subset 2\mathbb{N}$.
@@
Each orientation-preserving homeomorphism $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ induces the linear map $H_{1}(h)$ that preserves the intersection form on $H_{1}(S_{\mathrm{g}};\mathbb{Z})\cong\mathbb{Z}^{2\mathrm{g}}=\left<a_{1},\ldots,a_{\mathrm{g}},b_{1},\ldots,b_{\mathrm{g}}\right>$:
@@
Suppose $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ is an orientation-reversing homeomorphism. Thus it induces $-\rm{id}$ on $H_{2}(S_{\mathrm{g}})={\mathbb{Z}}$ and so the matrix $A$ of $H_{1}(h)$ in the standard basis satisfies
@@
In particular, one can show that $\det(A)=(-1)^{\mathrm{g}}$. The matrix $M=I_{\mathrm{g}}\oplus(-I_{\mathrm{g}})$ is clearly antisymplectic and $\det(M)=(-1)^{\mathrm{g}}$. If $A$ is antisymplectic, then $AM$ is symplectic, and so $\det(AM)=1$. Therefore $\det(A)=(\det(M))^{-1}=(-1)^{\mathrm{g}}$.
It is not difficult to verify that for a square matrix $A$ of dimension $n$,
@@
Theorem 3.2.
Let $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ be an orientation-reversing quasi-unipotent homeomorphism (e.g. Morse–Smale diffeomorphism). Then $L(h^{odd})=0$ and so ${\mathcal{A}P}(h)\subset 2{\mathbb{N}}$, i.e. $h$ has no odd algebraic periods.
@@
For $n$ odd, $h^{n}$ is also orientation-reversing, so $L(h^{n})=\mathrm{t}r(H_{1}(h^{n}))$ and $H_{1}(h^{n})$ is antisymplectic. Therefore it suffices to show that $L(h)=0$ just for $h$.
The characteristic polynomial $p(x)$ of $H_{1}(h)$ satisfies (8), so if $p(x)=(x-\xi)^{k}q(x)$. then
@@
Thus if $\xi$ is a root of $p(x)$ of multiplicity $k$, then $p(x)$ has a root $-1/\xi$ of multiplicity $k$. In particular, if $\xi$ is a primitive root of unity of odd order $l$, then $-1/\xi$ is a primitive root of unity of order $2l$, and conversely, if $\xi$ is primitive of order $2l$, $l$ odd, then $\xi^{l}=-1$, so $-1/\xi$ is primitive of order $l$.
Therefore, $p(x)$ has a factor $\varphi_{l}(x)$, $l$ odd, with multiplicity $k$ if and only if it has a factor $\varphi_{2l}(x)$ with multiplicity $k$. Since the sum of the roots of $\varphi_{m}(x)$ is equal to $\mu(m)$, $L(h)$ is a sum of the summands: $\mu(l)+\mu(2l)=\mu(l)-\mu(l)=0$ for each odd $l$ such that $\varphi_{l}(x)$ is a factor of $p(x)$ and $\mu(4m)=0$ for each factor $\varphi_{4m}(x)$ of $p(x)$. Thus $L(h)=0$.
@@
The statement of Theorem 1.4 follows from Theorems 2.3, 3.1, and Theorem 3.2. ∎
@@
Note that all the statements related to genus $\mathrm{g}$ in Theorem 3.1 remain valid for Morse–Smale diffeomorphisms in Theorem 1.4.
@@
In fact Llibre and coauthors considered orientation-reversing case in [Ll-Si1] (respectively, orientation-preserving case in papers cited in [Ll-Si3]) and they did not take into account the topological restrictions that come from the structure of cohomological ring which are encoded in the formula (7) (or respectively in (6)), which resulted in these papers in listing also non-realizable algebraic periods. In particular, as we showed above, these restrictions force that there are no odd numbers in the set of algebraic periods for orientation-reversing homeomorphisms.

### v2

@@ 3. Proof of main theorem
In the following considerations we will use the notion of periodic expansion of an arithmetic function $\psi:{\mathbb{N}}\to{\mathbb{C}}$ introduced in [19], exposed widely in [14], and used in [10], [11]. In this language, an arithmetic function $\psi$ is represented as a series of elementary periodic functions
@@
where the coefficients $a_{k}(\psi)\in{\mathbb{C}}$ are given by the Möbius inversion formula applied to the sequence $(\psi(n))$ analogously as it is in (1) for the sequence $(L(f^{n}))$, and
@@
is the sum of $n$-powers of all roots of unity of degree $k$ and is called the elementary periodic function. Conceptually, the periodic expansion is a discrete correspondent of the Fourier expansion, arithmetically it is linked with the Ramanujan sums, and its coefficients are used to study the behavior of the number sequences which are related to the periodic points (see [6]). In our context, it is just a convenient language to handle the sequence we are studying.
@@
    There exist a closed orientable surface $S_{\mathrm{g}}$ and an orientation-preserving homeomorphism $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ and the genus $\mathrm{g}$ of $S_{\mathrm{g}}$ is equal to
@@
    There exist a closed orientable surface $S_{\mathrm{g}}$ and an orientation-reversing homeomorphism $f\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ if and only if ${\mathcal{A}}\subset 2{\mathbb{N}}$. Moreover, the genus $\mathrm{g}$ of $S_{\mathrm{g}}$ is equal to
@@
    There exist a closed non-orientable surface $N_{\mathrm{g}}$ and a homeomorphism $f\colon N_{\mathrm{g}}\to N_{\mathrm{g}}$ of algebraically finite type such that ${\mathcal{A}P}(f)={\mathcal{A}}$ and the genus $\mathrm{g}$ of $N_{\mathrm{g}}$ is equal to
@@
First, let us describe the following construction. Having a finite set ${\mathcal{A}}^{\prime}\subset{\mathbb{N}}$ and an assignment $\tau\colon{\mathcal{A}}^{\prime}\to{\mathbb{N}}$, for each $n\in{\mathcal{A}}^{\prime}$ consider a surface $\Sigma_{n}$ which is an orientable surface $S_{\tau(n),2}$ in the orientable case or a non-orientable surface $N_{\tau(n),2}$ in the non-orientable case. On each $\Sigma_{n}$, we take a periodic homeomorphism $f_{n}$ of order $\tau(n)$ which cyclically permutes its $1$-handles and is a rotation on each boundary component. More specifically, for an illustration, start with a cylinder $S^{1}\times[-1,1]$ with a rotation by $1/\tau(n)$ of a full angle on each circle $S^{1}\times\{t\}$, and remove $\tau(n)$ open disjoint discs $D_{1},\ldots,D_{\tau(n)}$ in such a way that the rotation cyclically permutes the discs. Now, glue $\tau(n)$ copies of $S_{1,1}$ (or $N_{1,1}$) along permuted boundary components and extend the rotation to a periodic homeomorphism of order $\tau(n)$ that cyclically permutes the attached handles.
Next, having surfaces $\Sigma_{n_{1}},\ldots,\Sigma_{n_{k}}$, let us join $\Sigma_{n_{i}}$ and $\Sigma_{n_{i+1}}$ for $11\leq i\leq k-$ 1 by identifying their boundary components with end circles of a cylinder $C_{i}=S^{1}\times[-1,1]$. By the construction, we get a surface $S$ with two boundary components and define $f$ to be equal to $f_{n_{i}}$ on $\Sigma_{n_{i}}$, and on each joining cylinder it is given by a homotopy between two rotations on end circles (see Figure 1).
@@
Figure 1. Orientation-preserving case.
(1) In the case of an orientable surface and an orientation-preserving homeomorphism, take ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{1\}$ if $1\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{1\}$ if $1\in{\mathcal{A}}$, with an assignment $\tau(n)=n$ for $n\in{\mathcal{A}}^{\prime}$. The above procedure provides a surface $S$ with an orientation-preserving homeomorphism $f$ (see Figure 1). Let us attach two discs to $S$ to get a closed orientable surface $\widetilde{S}$, and $f$ extends on $\widetilde{S}$, being an appropriate rotation on each of two discs. Taking one circle $S^{1}\times\{0\}$ for each cylinder $C_{i}$, we get a system of simple closed curves preserved by $f$ such that $f$ is periodic on the complement of its open tubular neighborhood $\bigcup{\rm Int}C_{i}$, so $f$ is of algebraically finite type. Moreover, let $B_{n}=\{a^{n}_{1},\ldots,a^{n}_{n},b^{n}_{1},\ldots,b^{n}_{n}\}$ consist of circles in $S_{n}$ such that $\bigcup_{n\in{\mathcal{A}}^{\prime}}B_{n}$ forms the standard (symplectic) basis for $\widetilde{S}$. Then, the matrix of $H_{1}(f)$ on $H_{1}(\widetilde{S})$ is a block diagonal matrix whose blocks correspond to $B_{n}$’s, and since $f(a^{n}_{j})=f_{n}(a^{n}_{j})=a^{n}_{(j+1)\mod n}$ and similarly $f(b^{n}_{j})=b^{n}_{(j+1)\mod n}$, the $n$th block is the direct sum of two permutation matrices for a cycle of length $n$:
@@
and so it is not difficult to observe by definition of ${\rm reg}_{k}$ that $\mathrm{t}r(H_{1}(f^{l}))=\sum_{n\in{\mathcal{A}}^{\prime}}\linebreak 2{\rm reg}_{n}(l)$. By the definition of ${\mathcal{A}}^{\prime}$,
@@
where $a_{n}(f)=-2$ for $n\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$, and $a_{n}(f)=0$ for $n\notin{\mathcal{A}}$. Thus, ${\mathcal{A}P}(f)={\mathcal{A}}$.
(2) For the orientation-reversing case, we will show later that there are no odd algebraic periods (see Theorem 3.2). Assume that ${\mathcal{A}}\subset 2{\mathbb{N}}$ and take ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{2\}$ if $2\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{2\}$ if $2\in{\mathcal{A}}$ with an assignment $\tau(n)=n$ if $4$ divides $n$ and $\tau(n)=n/2$ otherwise. Again, the procedure provides an orientable surface $S$, and we attach a disc to one of its two boundary components (also naming the resulting surface $S$). We take its double $D(S)$, i.e. consider a copy $S^{\prime}$ of $S$ consisting of surfaces $\Sigma^{\prime}_{n}$ and cylinders $C^{\prime}_{i}$, and join $S$ and $S^{\prime}$ by a new cylinder $C=S^{1}\times[-1,1]$ along their boundary components. We have defined $f$ on $S$ and $S^{\prime}$, and on the new cylinder, let $f_{|S^{1}\times\{t\}}$ be the same as $f_{|{S^{1}\times\{\pm 1\}}}$.
Now, let $j\colon D(S)\to D(S)$ be orientation-reversing involution mapping points of $S$ into corresponding points of $S^{\prime}$, e.g., it can be seen as the reflection through the plane intersecting $D(S)$ in $S^{1}\times\{0\}\subset C$ and separating symmetrically $S$ and $S^{\prime}$, see Figure 2. The desired self-homeomorphism of $D(S)$ is $f^{\prime}=f\circ j$. Indeed, cutting $D(S)$ along circles from connecting cylinders, we get the decomposition into surfaces $\Sigma_{n}$ and $\Sigma^{\prime}_{n}$ such that $f^{\prime}$ maps $\Sigma_{n}$ homeomorphically onto $\Sigma^{\prime}_{n}$ and conversely. Thus, $(f^{\prime})^{2}$ is a self homeomorphism of each piece $\Sigma_{n}$ or $\Sigma^{\prime}_{n}$ on which one can check it is $n/2$-periodic. Hence, $f^{\prime}$ is of algebraically finite type.
[…]
Figure 2. Orientation-reversing case.
Moreover, it is clear that $(f^{\prime})^{2}(a^{n}_{j})=a^{n}_{(j+2)\mod\tau(n)}$, $(f^{\prime})^{2}(b^{n}_{j})=b^{n}_{(j+2)\mod\tau(n)}$ on $\Sigma_{n}$ and similarly for $\Sigma^{\prime}_{n}$. If $4$ divides $n$, then $\tau(n)=n$, and so $(f^{\prime})^{2}_{|\Sigma_{n}}$ has order $n/2$, thus the $n$th block of the matrix of $H_{1}(f^{\prime})$ is the direct sum of four permutation matrices for a cycle of length $n$. However, if $n$ is not divisible by $4$, then $\tau(n)=n/2$ is odd and $(f^{\prime})^{2}_{|\Sigma_{n}}$ still has order $\tau(n)=n/2$. Therefore, the characteristic polynomial of $f^{\prime}$ is equal to
@@
(3) Analogously, we construct a desired self-homeomorphism of a non-orientable surface. If ${\mathcal{A}}=\{1\}$, then just take ${\rm id}_{{\mathbb{R}}\!P^{2}}$. Otherwise, consider ${\mathcal{A}}^{\prime}={\mathcal{A}}\cup\{1\}$ if $1\notin{\mathcal{A}}$ or ${\mathcal{A}}^{\prime}={\mathcal{A}}\setminus\{1\}$ if $1\in{\mathcal{A}}$ with an assignment $\tau(n)=n$ for $n\neq 1$ and $\tau(1)=2$ if $1\in{\mathcal{A}}^{\prime}$. Moreover, in the preliminary construction for the non-orientable case, if $1\in{\mathcal{A}}^{\prime}$, we take $f_{1}={\rm id}$ on $\Sigma_{1}$ which is of genus $\tau(1)=2$. We get a non-orientable surface $S$ with two boundary components to which we attach discs as previously. Similarly as in (1), we have a self-homeomorphism $f$ of $S$ of algebraically finite type, which is $n$-periodic on each $\Sigma_{n}$, except $n=1$.
Let $B_{n}=\{a^{n}_{1},\ldots,a^{n}_{\tau(n)}\}$ be a set of circles in $\Sigma_{n}$ which are cyclically permuted by $f$ and such that elements of $\bigcup B_{n}$ represent standard generators of $H_{1}(S)\cong{\mathbb{Z}}^{g-1}\oplus{\mathbb{Z}}_{/2}$ with the relation $02\sum_{i,n}a^{n}_{i}=$ 0, where $\mathrm{g}$ is the genus of $S$. The trace of $H_{1}(f)$ is computed as the trace of $\widetilde{H_{1}(f)}$ on $H_{1}(S)/{\rm T}\cong{\mathbb{Z}}^{\mathrm{g}-1}$, which is induced by $H_{1}(f)$ after dividing by the torsion part ${\rm T}$ of $H_{1}(S)$. Thus, $H_{1}(S)/{\rm T}\cong\left(\bigoplus_{i,n}{\mathbb{Z}}a^{n}_{i}\right)/\left<\sum_{i,n}a^{n}_{i}\right>$ has a basis $\left(\bigcup B_{n}\right)\setminus\{a\}$ for some fixed $a\in B_{n_{0}}$, and $a$ is represented as the vector $(-1,\ldots,-1)$ in this basis. Hence, the matrix of $\widetilde{H_{1}(f)}$ is a block matrix if $n_{0}=1\in{\mathcal{A}}^{\prime}$, or almost a block matrix if $n_{0}\neq 1$, i.e. the structure of a block matrix is disturbed only in the column which corresponds to a basis element $a^{n_{0}}_{i}$ such that $f(a^{n_{0}}_{i})=a$ (if we order the basis in such a way that $a^{n_{0}}_{i}$ is the last element, then the matrix is upper block triangular). However, in the second case, the characteristic polynomial of $\widetilde{H_{1}(f)}$ is still the product of characteristic polynomials of permutation matrices for cycles of length $n$ for $n\neq n_{0}$, the identity matrix of dimension $2$ if $1\in{\mathcal{A}}^{\prime}$, and the characteristic polynomial of
@@
which is the companion matrix of the polynomial $x^{n_{0}-1}+\ldots x+1=(x^{n_{0}}-1)/(x-1)$. Since the trace of the companion matrix of a polynomial is equal to the sum of its roots, by formula (3), the trace of the $l$-power of the matrix (5) is equal to ${\rm reg}_{n_{0}}(l)-{\rm reg}_{1}(l)$. Thus, in both the cases we get
@@
where $a_{n}(f)=-1$ for $n\in{\mathcal{A}}\setminus\{1\}$, $a_{1}(f)=2$ if $1\in{\mathcal{A}}$, and $a_{n}(f)=0$ for $n\notin{\mathcal{A}}$. Hence, ${\mathcal{A}P}(f)={\mathcal{A}}$. ∎
To complete the proof, we are left with showing that ${\mathcal{A}P}(f)\cap(2{\mathbb{N}}-1)=\emptyset$ if $f$ is orientation-reversing because we have shown only the realization for ${\mathcal{A}P}(f)\subset 2{\mathbb{N}}$.
@@
Each orientation-preserving homeomorphism $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ induces the linear map $H_{1}(h)$ that preserves the intersection form on $H_{1}(S_{\mathrm{g}};\mathbb{Z})\cong\mathbb{Z}^{2\mathrm{g}}=\left<a_{1},\ldots,a_{\mathrm{g}},b_{1},\ldots,b_{\mathrm{g}}\right>$,
@@
Suppose $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ is an orientation-reversing homeomorphism. Thus, it induces $-\rm{id}$ on $H_{2}(S_{\mathrm{g}})={\mathbb{Z}}$, and so the matrix $A$ of $H_{1}(h)$ in the standard basis satisfies
@@
In particular, one can show that $\det(A)=(-1)^{\mathrm{g}}$. The matrix $M=I_{\mathrm{g}}\oplus(-I_{\mathrm{g}})$ is clearly antisymplectic, and $\det(M)=(-1)^{\mathrm{g}}$. If $A$ is antisymplectic, then $AM$ is symplectic, and so $\det(AM)=1$. Therefore, $\det(A)=(\det(M))^{-1}=(-1)^{\mathrm{g}}$.
It is not difficult to verify that for a square matrix $A$ of dimension $n$, its characteristic polynomial $\chi_{A}(x)$ satisfies
@@
Theorem 3.2 (cf. [3]).
Let $h\colon S_{\mathrm{g}}\to S_{\mathrm{g}}$ be an orientation-reversing quasi-unipotent homeomorphism (e.g. Morse–Smale diffeomorphism). Then, $L(h^{odd})=0$, and so ${\mathcal{A}P}(h)\subset 2{\mathbb{N}}$, i.e. $h$ has no odd algebraic periods.
@@
For $n$ odd, $h^{n}$ is also orientation-reversing, so $L(h^{n})=\mathrm{t}r(H_{1}(h^{n}))$ and $H_{1}(h^{n})$ is antisymplectic. Therefore, it suffices to show that $L(h)=0$ just for $h$.
The characteristic polynomial $p(x)$ of $H_{1}(h)$ satisfies (8), so if $p(x)=(x-\xi)^{k}q(x)$, then
@@
Thus, if $\xi$ is a root of $p(x)$ of multiplicity $k$, then $p(x)$ has a root $-1/\xi$ of multiplicity $k$. In particular, if $\xi$ is a primitive root of unity of odd order $l$, then $-1/\xi$ is a primitive root of unity of order $2l$, and conversely, if $\xi$ is primitive of order $2l$, $l$ odd, then $\xi^{l}=-1$, so $-1/\xi$ is primitive of order $l$.
Let $\varphi_{l}(x)$ be the $l$th cyclotomic polynomial. Then, $p(x)$ has a factor $\varphi_{l}(x)$, $l$ odd, with multiplicity $k$ if and only if it has a factor $\varphi_{2l}(x)$ with multiplicity $k$. Since the sum of the roots of $\varphi_{m}(x)$ is equal to $\mu(m)$, $L(h)$ is a sum of the summands: $\mu(l)+\mu(2l)=\mu(l)-\mu(l)=0$ for each odd $l$ such that $\varphi_{l}(x)$ is a factor of $p(x)$, and $\mu(4m)=0$ for each factor $\varphi_{4m}(x)$ of $p(x)$. Thus, $L(h)=0$.
@@
The statement of Theorem 1.3 follows from Theorems 2.3, 3.1, and 3.2. ∎
@@
Note that all the statements related to the genus $\mathrm{g}$ in Theorem 3.1 remain valid for Morse–Smale diffeomorphisms in Theorem 1.3.
@@
In fact, Llibre and coauthors considered the orientation-reversing case in [16] (respectively, orientation-preserving case in papers cited in [18]), and they did not take into account the topological restrictions that come from the structure of the cohomological ring which are encoded in condition (7) (or respectively in (6)), which resulted in these papers in listing also non-realizable algebraic periods. In particular, as we showed above, these restrictions force that there are no odd numbers in the set of algebraic periods for orientation-reversing homeomorphisms.